News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Tuesday 21 September 1999 9.00am EST

During the last 24 hours a series of GTO observations with both HRC and ACIS were completed, including the targets Abell 426 and M82. The spacecraft systems were nominal during this period.

Analysis of ACIS data from the last 2 radiation passages indicates no further degradation in the front-side chip energy resolution. The back side-chip data continues to show no change since launch. The initial results from the diagnostic tests run over the weekend show an increasing trend of energy resolution loss as the focal plane operated at higher temperatures (-110 to -60 deg C range), then indication of turn over at -60 deg C. These results are expected to help identify the type of effect causing the degradation in energy resolution.

The next load was prepared yesterday and reviewed overnight, and will perform a set of ACIS GTO observations including 30-Doradus and Abell 3667. Additional calibration measurements will also be made with Capella. ACIS internal calibration measurements will be taken prior to the next radiation belt entry on Tuesday evening, and prior to the next eclipse on Wedensday at 5am.





Return to Status Reports